# Shavette: Safe Undervolting of Neural Network Accelerators via Algorithm-level Error Detection

### v2

@@ 6. Conclusion
This study demonstrates the existence of a significant voltage margin in the default settings of a commercial GPU and possibility of reliably removing it. In computations of DNN models, the margin can be safely optimized by incorporating ABFT schemes. The time and energy overheads from ABFT are low, e.g., around 3.5% and 1.1% respectively. Nearly a quarter of energy could be saved without sacrificing accuracy of the neural network. The results indicate a promising avenue for enhancing the energy efficiency of AI accelerators, without expensive HW techniques or having to compromise reliability.

### v3

@@ 6 Conclusion
In computations of DNN models, the voltage margins of the accelerator can be safely optimized as DNN models mostly comprised of operations in which ABFT schemes can be integrated into. This enables timely detection of voltage induced errors with minimal overheads. The time and energy overheads from ABFT are low, e.g., around 3.5% and 1.1% respectively. This work demonstrated that nearly a quarter of energy could be saved without sacrificing accuracy of the model. The results indicate a promising avenue for enhancing the energy efficiency of AI accelerators, without expensive HW techniques or having to compromise reliability.
Acknowledgment
This work was supported by 6G Flagship (Grant Number 369116) funded by the Research Council of Finland.